the sum of two numbers is 31. if the smaller is doubled and the larger is increased by 11, the resulting numbers will be equal. find both numbers.
Answer by mananth(16946)

x+y=31 x>y
x+11=2y
x-2y=-11
Multiply the first equation by 2
2x+2y=62
x-2y=-11
add them up
3x=51
x= 17
x+y=31
Therefore y=14
the numbers are 17 & 14